OpenStudy (anonymous):

What is the probability of getting heads when flipping a coin and getting a number greater than or equal to 5 when rolling a single die? 1. 1/3 2. 1/6 3. 1/12 4. 1/4

Directrix (directrix):

When tossing a coin, there are 2 outcomes: heads and tails. Probability of Heads = 1/2 because there are two possible outcomes and there is only one way to get Heads.

Directrix (directrix):

@vesper When you toss a single die, how many different outcomes are there? Recall that the die is cubical.

Directrix (directrix):

Yo, @vesper Say something. How many faces on a die?

OpenStudy (anonymous):

@Directrix 6

Directrix (directrix):

Good.

Directrix (directrix):

Now, we want to find the probability of this: number greater than or equal to 5 when rolling a single die

Directrix (directrix):

As you said, there are 6 faces on the die. They have "spots" on them: 1, 2, 3, 4, 5, or 6.

Directrix (directrix):

There are two ways to get a die with 5 spots or higher. That is to toss a 5 or to toss a 6. TWO ways.

Directrix (directrix):

So, the probability of getting a 5 or a 6 when tossing a single die is 2/6 because there are 6 possible outcomes and two of them are what we want.

Directrix (directrix):

Probability (Heads on coin AND 5 or 6 on the die) = 1/2 * 2/6

Directrix (directrix):

1/2 * 1/3 = ? @vesper Note: the 2/6 on the die simplifies to 1/3

OpenStudy (anonymous):

1/6 and yes got it that made perfect sense to me all thank you very much for that now i see where i was messing up cause i thought the answer was 1/3

Directrix (directrix):

Okay. Good to see that you found the error of your ways.

OpenStudy (anonymous):

@Directrix yes i did ty again for the helful walkthrough

Directrix (directrix):

Alrighty. You are welcome.